Number 7B - What is your dollar weighted rate of return? 7 – You manage a fund with an expected rate of return of 11% and a standard
Answer #1

1.
Expected return=80%*11%+20%*1%=9.00%

Standard deviation=80%*19%=15.20%

2.
Let w be the weight in fund
w*11%+(1-w)*1%=12%
=>w=(12%-1%)/(11%-1%)=1.1

% invested in fund=1.1*100%=110%

Standard deviation=1.1*19%=20.90%